# Foundation Cures Personalization: Improving Personalized Models’ Prompt Consistency via Hidden Foundation Knowledge

### v2

@@ 8.1 Triggering Hidden Foundation Knowledge in Personalization Baselines
The core architecture of FreeCure is based on a dual denoising paradigm of the FD and PD process, which triggers hidden foundation knowledge in personalization models. This paradigm deactivates the identity embeddings which are integrated into the cross-attention layers. In Table 4 we illustrate how we deactivate identity embedding for different baselines mentioned in the main paper since they have various identity embedding fusion strategies and prompt templates.
@@
Table 4: Strategies for implementing dual denoising paradigms on different baselines. “black curly hair” is used here as an example and can be replaced with other prompts containing different facial attributes.
FastComposer and Face-Diffuser share the same prompting template, wherein their identity embeddings are integrated into the “<P>” word, which is initialized with class-specific words (e.g., man, woman, boy, girl). To set FD process, we keep this “<P>” as its original plain text’s embedding without integrating any identity information. In the case of PuLID, which utilizes independent residual cross-attention adapters for identity fusion, we initialize its identity embedding as a zero tensor of the same dimensionality.
8.2 Attribute Segmentation Models
As mentioned in Section 5, we use two mainstream segmentation models: BiSeNet and Segment-Anything models for attribute-aware mask generation. BiSeNet  [^1] https://github.com/CoinCheung/BiSeNet is a popular framework for face parsing that is capable of generating semantic masks corresponding to facial attributes. Table 5 shows BiSeNet’s prediction label and corresponding facial attributes. BiSeNet can address the majority of facial attributes, thereby verifying the robustness of FreeCure. Additionally, FreeCure is designed to be seamlessly integrated with various segmentation models, not limited to BiSeNet. To validate this flexibility, we have showcased results that utilize Segment-Anything as the face parsing model in Section 9.3.
@@ 8.2 Attribute Segmentation Models
Table 5: Information for facial attribute and label relationship of BiSeNet. All critical attributes for the human face are included.
8.3 Details of Metrics
Prompt Consistency (PC, also known as CLIP-T). We leverage the official implementation of the vision transformer [^2] https://github.com/openai/CLIP provided by OpenAI. We utilize its image and text encoder to derive image and textual embeddings respectively. Ultimately, the cosine similarity between these two embeddings is computed as a quantitative measure of prompt consistency.
8.4 Facial Prompts for Evaluation
@@
a <S> with sunglasses and silver earrings
@@
9 More Results
9.1 Supplementary Comparison of Results for Main Paper
To better demonstrate the efficacy of FreeCure, Figure 12 presents a comparative analysis of the results generated by the FD, PD, and enhanced outputs by FreeCure linked to Figure 6. The results indicate that FreeCure faithfully transfers target prompt-aligned attributes from the FD into the PD processes while preserving the original identity information.
9.2 Results on Non-Celebrity References
@@
In Figure 13, we provide additional experimental results on non-celebrity reference images to further validate the consistent performance of FreeCure. This robust performance shows the potential of FreeCure to enhance various personalization models in practical applications, as the ability to handle non-celebrity personalization is a critical requirement in real-world scenarios.
9.3 Integration with Segment-Anything for Attribute Extraction
Figure 14 showcases FreeCure’s consistent performance when integrated with more advanced segmentation models such as Segment-Anything. For instance, the attribute mask cannot be handled by BiSeNet since its label is absent. However, by replacing BiSeNet with Segment-Anything and specifying the target prompt as “mask,” valid semantic masks can be accurately generated, thereby guaranteeing effective attribute enhancement through FASA. Given that Segment-Anything extracts attributes based on flexible prompts, its seamless integration significantly enhances the versatility of FreeCure, enabling it to address a broader range of scenarios.
10 More Analysis
10.1 More Comparison Results w/ and w/o Identical Initial Noise
We conduct extensive experiments to evaluate the impact of initial noise conditions. Figure 15 demonstrates that under identical initial noise conditions, FreeCure consistently enhances target attributes across different input reference images. This finding highlights the robustness and practical applicability of FreeCure. Figure 16 illustrates that FreeCure can successfully inject attribute information from the FD process to the PD process even if their final generated faces have various locations, angles, and sizes. This finding particularly demonstrates the robustness of the FASA module.
10.2 More Visualization of FASA
We provide a more detailed visualization of FASA in Figure 17, which substantiates the claim that FASA enhances facial attributes in a fine-grained manner. For instance, FASA effectively captures and faithfully transfers information of attributes localized in small areas (e.g. eyes, pearl earrings). Furthermore, in regions unrelated to the target attributes, FASA maintains a strong alignment with the original PD attention maps, demonstrating that it preserves the core functionality of the personalization models.
10.3 Ablations of Applying FASA at Different Positions in the U-Net
We conduct detailed ablation studies to determine the optimal placement of the FASA module within the U-Net, as illustrated in Figure 18. When FASA is applied exclusively to the downsampling and middle blocks, its effectiveness is limited. In contrast, applying FASA to the upsampling blocks yields the most significant performance improvements. Furthermore, adding FASA to the middle and downsampling blocks provides negligible additional benefits when the upsampling blocks are already utilized, as the upsampling blocks alone are sufficient to accurately transfer attribute information from the FD to the PD process. Based on these findings, we conclude that applying FASA to the U-Net’s upsampling blocks represents the optimal configuration.
10.4 More Studies on Identity Embedding
Figure 19 illustrates the impact of identity embedding interpolation when integrated into the cross-attention layers of different blocks, as mentioned in Section 3. When cross-attention maps of the FD process are injected into the downsampling blocks or middle blocks, the changes in output are minimal, even if all identity embedding’s cross-attention maps are replaced ($\alpha=1$). It is only when the interpolation is applied to the upsampling blocks that significant degradation of identity information is evident, while other facial attributes are effectively restored. In summary, the identity embedding exerts its most significant influence within the upsampling blocks of the U-Net. However, a small value of $\alpha$ can cause significant identity information loss, supporting the argument in Section 3, that the well-trained cross-attention layers for identity information extraction is susceptible.
@@ 10.4 More Studies on Identity Embedding
Figure 12: Corresponding comparative FD & PD outputs for results in Figure 6 in the main paper. (a) refers to original personalization outputs, (b) refers to foundation outputs, and (c) refers to outputs enhanced with proposed FreeCure.
Figure 13: More results for non-celebrity references.
@@
Figure 14: Results for FreeCure’s integration with Segment-Anything. FreeCure can seamlessly integrate with more advanced segmentation models such as Segment-Anything for extraction of a boarder range of attributes. Target prompt: “a person wearing a mask”
@@
Figure 19: Identity embedding interpolation’s influence on cross-attention layers of different blocks in the U-Net. It is applied on: downsampling blocks (top), bottleneck blocks (middle), and upsampling blocks (bottom).

### v3

@@ A.2.1 Triggering Hidden Foundation Knowledge in Personalization Baselines
The core architecture of FreeCure is based on a dual denoising paradigm of the FD and PD process, which triggers hidden foundation knowledge in personalization models. This paradigm deactivates the identity embeddings which are integrated into the cross-attention layers. In Table 4 we illustrate how we deactivate identity embedding for different baselines mentioned in the main paper since they have various identity embedding fusion strategies (e.g., LoRA, adapters) and prompt templates.
Table 4: Strategies for implementing dual denoising paradigms on different baselines. “black curly hair" is used here as an example and can be replaced with other facial attributes.
@@
InstantID a <P> with black curly hair a <P> image with black curly hair
@@
PuLID a <P> with black curly hair a <P> image with black curly hair
InfiniteYou a <P> with black curly hair a <P> image with black curly hair
FastComposer [64], Face-Diffuser [62] and PhotoMaker [34]share the similar prompting template, wherein their identity embeddings are integrated into the “<P>" word, which is initialized with class-specific words (e.g., man, woman, boy, girl). To establish the FD process, we maintain the original plain text embedding for “<P>" without incorporating any identity information, effectively equivalent to applying a null identity embedding. Face2Diffusion [49] integrated the identity embedding into the trigger word “f". Therefore, we replace the identity embedding with a zero-valued tensor before its merging process. For baselines such as InstantID [59], PuLID [18], and InfiniteYou [27] that employ residual cross-attention adapters for identity fusion, we initialize their identity embeddings as zero tensors with matching dimensionality while preserving the original textual embeddings unchanged.
A.2.2 Attribute Segmentation Models
As mentioned in Sec.5, we use two mainstream segmentation models: BiSeNet and Segment-Anything models for attribute-aware mask generation. BiSeNet  [^1] https://github.com/CoinCheung/BiSeNet is a popular framework for face parsing that is capable of generating semantic masks corresponding to facial attributes. Table 5 shows BiSeNet’s prediction label and corresponding facial attributes, demonstrating that it can address the majority of facial attributes, thereby verifying the robustness of FreeCure. Additionally, FreeCure is designed to be seamlessly integrated with various segmentation models, not limited to BiSeNet. We have showcased results that utilize Segment-Anything as the face parsing model in Sec.A.3.4.
Table 5: Information for facial attribute and label relationship of BiSeNet.
@@ A.2.2 Attribute Segmentation Models
a <S> with sunglasses, frowning worriedly
@@
A.2.3 Facial Prompts for Evaluation
A.2.4 Details of Metrics
Prompt Consistency (PC, also known as CLIP-T). We leverage the official implementation of the vision transformer [^2] https://github.com/openai/CLIP provided by OpenAI.
A.3 More Results
A.3.1 More Results on Celebrity & Non-Celebrity References
@@
In Fig.11, 12 and 13, we provide additional experimental results on more reference images to further validate the consistent performance of FreeCure. This robust performance shows the potential of FreeCure to enhance various personalization models in practical applications, as the ability to handle non-celebrity personalization is a critical requirement in real-world scenarios.
Figure 11: More results for baselines implemented with SDv1.5, including FastComposer, Face-Diffuser and Face2Diffusion. We provide personalization results including non-celebrity identities.
Figure 12: More results for baselines implemented with SDXL, including InstantID, PhotoMaker and PuLID-SDXL. We provide personalization results including non-celebrity identities.
Figure 13: More results for baselines implemented with FLUX, including PuLID-FLUX and InfiniteYou. We provide personalization results including non-celebrity identities.
A.3.2 Supplementary Comparison of Results for Main Paper
To better demonstrate the efficacy of FreeCure, Fig.14, 15 present a comparative analysis of the results generated by the FD, PD, and enhanced outputs by FreeCure linked to Fig.5. The results indicate that FreeCure faithfully transfers target prompt-aligned attributes from the FD into the PD processes while preserving the original identity information.
@@ A.3.2 Supplementary Comparison of Results for Main Paper
Figure 14: Corresponding comparative FD & PD outputs for results in the main submission. (FastComposer, Face-Diffuser and FaceDiffusion) (a) refers to original personalization outputs, (b) refers to foundation outputs, and (c) refers to outputs enhanced with proposed FreeCure.
Figure 15: Corresponding comparative FD & PD outputs for results in the main paper. (InsantID, PhotoMaker and PuLID) (a) refers to original personalization outputs, (b) refers to foundation outputs, and (c) refers to outputs enhanced with proposed FreeCure.
A.3.3 More Detailed Quantitative Results for Prompt Consistency with Multiple Attributes.
As supplementary results for Table.2, we provide source prompt consistency results of each baseline considering prompts with various attributes in Table.7. It shows that as the prompt becomes increasingly complicated, all baselines are facing more decreasing in prompt consistency, which can limit their application in real scenarios. The results also support the fact that FreeCure’s improvement in PC becomes more significant as the complexity of the prompt increases.
FastComposer 18.83 18.05 16.92
Face-Diffuser 21.68 20.98 18.02
Face2Diffusion 22.54 21.98 20.54
InstantID 22.43 21.81 20.95
InstantID + FreeCure 23.98 (+6.91%) 23.52 (+7.84%) 23.11 (+10.31%)
PhotoMaker 23.51 23.01 22.57
PhotoMaker + FreeCure 25.04 (+6.51%) 24.91 (+8.26%) 24.64 (+11.19%)
PuLID (SDXL) 25.56 25.12 24.43
PuLID (SDXL) + FreeCure 26.30 (+3.34%) 25.96 (+5.32%) 25.73 (+3.55%)
PuLID (FLUX) 23.29 22.08 21.32
PuLID (FLUX) + FreeCure 25.52 (+9.57%) 24.49 (+10.91%) 23.87 (+11.96%)
InfiniteYou 25.01 23.19 22.43
InfiniteYou + FreeCure 25.98 (+3.88%) 24.92 (+7.46%) 24.45 (+9.01%)
A.3.4 Integration with Segment-Anything for Attribute Extraction
Fig.16 showcases FreeCure’s consistent performance when integrated with more advanced segmentation models such as Segment-Anything. For instance, the attribute mask cannot be handled by BiSeNet since its label is absent. However, by replacing BiSeNet with Segment-Anything and specifying the target prompt as “mask," valid semantic masks can be accurately generated, thereby guaranteeing effective attribute enhancement through FASA. Given that Segment-Anything extracts attributes based on flexible prompts, its seamless integration significantly enhances the versatility of FreeCure, enabling it to address a broader range of scenarios.
@@ A.3.4 Integration with Segment-Anything for Attribute Extraction
Figure 16: Results for FreeCure’s integration with Segment-Anything. FreeCure can seamlessly integrate with more advanced segmentation models such as Segment-Anything for extraction of a boarder range of attributes. Target prompt: “a person wearing a mask".
A.3.5 Analysis on non-facial attributes enhancement
We have conducted supplementary experiments about attributes beyond facial elements. Following the prompt set of PuLID, we evaluated prompts involving common objects, clothing, and other typical non-facial attributes, using Segment Anything for mask generation. The experimental results across several baselines are shown in Table.8:
Table 8: Quantitative comparison of on non-facial attributes.
PC(%) $\uparrow$ IF(%) $\uparrow$ Face Div. (%) $\uparrow$
Face2Diffusion 22.87 41.08 41.34
Face2Diffusion + FreeCure 23.34 (+2.06%) 40.75 (-0.80%) 41.91 (+1.38%)
InstantID 22.43 65.55 47.73
InstantID + FreeCure 23.51 (+4.81%) 65.34 (-0.32%) 48.06 (+0.69%)
PhotoMaker 24.67 52.14 46.02
PhotoMaker + FreeCure 25.37 (+2.84%) 51.77 (-0.71%) 46.58 (+1.21%)
PuLID (SDXL) 27.43 59.41 41.97
PuLID (SDXL) + FreeCure 28.13 (+2.55%) 59.03 (-0.64%) 42.05 (+0.19%)
Experimental results indicate that the facial personalization model exhibits a weaker copy-paste effect in non-facial regions, leading to less pronounced improvements in prompt consistency compared to the facial attributes specifically addressed in our submission. However, we observed that FreeCure still demonstrates significant restoration effects for attributes such as headphones and hats, which are in close proximity to the face (see Fig.17). Moreover, since the restoration occurs outside the facial region, FreeCure has a more negligible negative impact on identity fidelity.
Figure 17: Results on non-facial attributes. FreeCure can still improve prompt consistency on non-facial attributes, especially on objects whose location is close to face.
A.4 More Analysis
A.4.1 Non-interference Enhancement of Multiple Facial Attributes
We validate that the subsequent enhancement processes of FreeCure do not impact attributes that have already been enhanced. Fig.18 visualizes the intermediate results at FreeCure’s different stages. When APG is used in the latter stages, attributes previously enhanced through FASA remain unaffected. This illustrates that FreeCure’s improvement across multiple attributes is highly robust and consistent.
Figure 18: Demonstration of FreeCure’s non-interference manner. top row: personalization models’ outputs; middle row: intermediate enhanced results with FASA; bottom row: final results after APG enhancement.
A.4.2 More Robustness Justification under Inaccurate Attribute Masks
Although FreeCure rely on attribute segmentation process based on pretrained semantic segmentation models. Its performance does not sensitive to the accuracy of these models. We design two-way validation experiments that introduce inaccuracy to attribute masks. First, we apply a dilation operation on masks, which might include area which does not belong to the target attribute area. Second, we apply an erosion operation, which might cause loss of some useful information about attributes. To create significant distortion of masks, we set the receptive field of both dilation and erosion to 25 pixels. We conduct experiments on several baselines, and the Table.9 below reports the results for respective prompt consistency (PC) and identity fidelity (IF).
Table 9: PC and IF Performance for Inaccurate Masks
original IF original PC dialated IF dialated PC eroded IF eroded PC
InstantID 62.01 23.62 62.23 23.45 62.23 23.21
PhotoMaker 50.15 24.91 50.01 24.87 50.44 24.34
PuLID-SDXL 56.95 26.05 56.74 25.51 57.01 25.36
PuLID-FLUX 72.61 24.78 72.39 24.52 72.97 24.32
The results demonstrate that identity fidelity (IF) remains largely consistent across both dilated and eroded mask conditions, with slight improvements observed in eroded masks. This occurs because mask erosion preserves certain attributes similar to the reference image. As for prompt consistency, both dilation and erosion cause minimal degradation, though erosion exhibits marginally greater impact. We consider that this stems from cases where fine-grained attribute restoration (e.g., iris color, earrings) may fail when the mask completely degenerates due to erosion.
Generally, the performance degradation of FreeCure due to inaccurate masks is limited. Moreover, in practical applications, specialized segmentation models (e.g., BiSeNet, Segment Anything) exhibit high robustness, making failure cases caused by mask inaccuracies even rarer than observed in these deliberately designed experiments. This further validates the inherent robustness of the FreeCure framework.
A.4.3 More Comparison Results w/ and w/o Identical Initial Noise
We conduct extensive experiments to evaluate the impact of initial noise conditions. Fig.19 demonstrates that under identical initial noise conditions, FreeCure consistently enhances target attributes across different input reference images. This finding highlights the robustness and practical applicability of FreeCure. Fig.20 illustrates that FreeCure can successfully inject attribute information from the FD process to the PD process even if their final generated faces have various locations, angles, and sizes. This finding particularly demonstrates the robustness of the FASA module.
@@ A.4.3 More Comparison Results w/ and w/o Identical Initial Noise
A.4.4 More Detailed Quantitative Results for Ablation Studies
Table.10 provides a supplementary details for ablation studies mentioned in Table.3. The results indicate that both FASA and APG consistently improve prompt consistency across all original personalization models. Notably, FASA yields more substantial improvements compared to APG, aligning with our findings in Sec.5.3.
Table 10: Quantitative ablation analysis for FASA and APG’s independent effect on overall performance. The table provides detailed ablation studies on each baselines.
FastComposer 20.91 41.94 27.91
19.03 42.84 26.35
Face-Diffuser 21.86 57.92 31.74
20.74 58.11 30.60
Face2Diffusion 22.73 39.30 28.80
21.98 39.47 28.24
InstantID 23.02 62.45 33.64
22.76 62.51 33.40
PhotoMaker 24.76 50.62 33.25
23.61 50.86 32.25
PuLID (SDXL) 25.86 57.85 35.74
25.54 57.24 35.33
PuLID (FLUX) 24.12 72.98 36.26
23.86 73.03 35.97
InfiniteYou 24.87 78.34 37.75
24.32 78.67 37.15
A.4.5 Ablations of Applying FASA at Different Positions in the Denoising Model
We conduct detailed ablation studies to determine the optimal placement of the FASA module within the denoising model, as illustrated in Fig.21. When FASA is applied exclusively to the downsampling and middle blocks, its effectiveness is limited. In contrast, applying FASA to the upsampling blocks yields the most significant performance improvements. Furthermore, adding FASA to the middle and downsampling blocks provides negligible additional benefits when the upsampling blocks are already utilized, as the upsampling blocks alone are sufficient to accurately transfer attribute information from the FD to the PD process. Based on these findings, we conclude that applying FASA to the denoising model’s upsampling blocks represents the optimal configuration.
A.4.6 More Visualization of FASA
We provide a more detailed visualization of FASA in Fig.22, which substantiates the claim of Sec. 5.2 that FASA enhances facial attributes in a fine-grained manner. For instance, FASA effectively captures and faithfully transfers information of attributes localized in small areas (e.g. eyes, pearl earrings). Furthermore, in regions unrelated to the target attributes, FASA maintains a strong alignment with the original PD attention maps, demonstrating that it preserves the core functionality of the personalization models.
@@ A.4.6 More Visualization of FASA
A.4.7 More Studies on Identity Embedding
Fig.23 illustrates the impact of identity embedding interpolation when integrated into the cross-attention layers of different blocks, as mentioned in Sec.3. When cross-attention maps of the FD process are injected into the downsampling blocks or middle blocks, the changes in output are minimal, even if all identity embedding’s cross-attention maps are replaced ($\alpha=1$). It is only when the interpolation is applied to the upsampling blocks that significant degradation of identity information is evident, while other facial attributes are effectively restored. In summary, the identity embedding exerts its most significant influence within the upsampling blocks of the denoising model. However, a small value of $\alpha$ can cause significant identity information loss, supporting the argument in Sec.3, that the well-trained cross-attention layers for identity information extraction is susceptible.
Figure 23: Identity embedding interpolation’s influence on cross-attention layers of different blocks in the denoising model. It is applied on: downsampling blocks (top), bottleneck blocks (middle), and upsampling blocks (bottom).
We have conducted a detailed analysis using finer-grained $\alpha$ values and their corresponding quantitative outcomes. Below are the results:
Table 11: Fined-grained Quantitative Metrics for Examples of Cross-attention Interpolation in Sec.3
$\alpha$ 0 0.2 0.4 0.6 0.8 1
top Identity Fidelity 50.23 24.53 10.34 6.42 3.43 2.1
Prompt Consistency 24.98 25.5 26.39 28.54 30.43 33.01
bottom Identity Fidelity 52.34 49.52 29.12 15.75 6.32 3.2
Prompt Consistency 26.76 30.34 31.09 32.98 34.23 35.02
The results demonstrate that identity fidelity declines sharply even when only a small portion of the original cross-attention map is altered, suggesting that identifying a well-optimized sweet spot through cross-attention manipulation alone is highly challenging. In summary, difficulties in cross-attention manipulation provide us with a more solid motivation in the exploration of self-attention modules.
A.4.8 Inference Time Analysis
We conduct experiments on a single H800 GPU with 80 GB VRAM to measure each baseline’s inference time before and after applying FreeCure. Table.12 below reports the results (“*" means 30-step denoising due to official guidance and 50-step denoising is applied for the rest baselines)
Table 12: Runtime analysis
Baseline (seconds) Baseline + FreeCure (seconds)
Stage 1 Stage 2 with FASA APG Total
FastComposer 1.79 1.96 2.19 1.83 5.98
Face-Diffuser 2.13 2.63 2.33 1.92 6.88
Face2Diffusion* 1.13 1.25 1.47 1.23 3.95
PhotoMaker 6.52 7.04 10.03 6.1 23.17
InstantID* 7.12 7.63 10.65 6.23 24.51
PuLID-SDXL 7.43 8.02 10.96 7.28 26.26
PuLID-FLUX* 12.84 14.89 16.86 12.95 44.7
InfineteYou* 8.34 10.23 13.53 9.03 32.79
It is true that introducing extra calculation can lead to longer runtime, we would like to emphasize that most training-free methods introduce extra processes, including inversion operation, denoising processes, and extra attention calculation. Plus, comparing to the fact that most encoder-based personalization methods require a long range time of data collection, preprocessing and tuning (this may consume a large array of GPUs and thousands of GPU-hours), our training-free method provides an innovative perspective for prompt consistency improvement by the knowledge in themselves, without the need for designing new model architecture and dataset curation.
A.4.9 User Study Template
Fig.24 show an example of our proposed user study on the performance of FreeCure.
Figure 24: An example of user study. We use scoring strategy to collect user assessments about the samples with/without FreeCure.